

********************************************************************
UARS WEEKLY STATUS REPORT - 22 Oct 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 61% of
each day's data amounting to nearly 685 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004179 has been reprocessed and archived at the GSFC DAAC. HEPSA Level 2
data files from 2004180 to 2004189 are processed and waiting for archiving.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004139.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 October 2004 (Orbit 71571, GMT Day 283) through
15 October 2004 (Orbit 71676, GMT Day 289).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 284/1248 71594 PEM Z/N DAY ONLY TO FULL
PEM 284/1354 71595 PEM AXIS 1 & 2 ON
PEM 286/0043 71616 POWER ON (OFF TO FULL)
PEM 286/1607 71626 PEM RESET `TIMING ERROR'
The beta angle ranged between the angles of 61.3 degrees to 65.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 02 November 2004 (DOY 307).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 0.94 degrees C. Additional
battery performance data for each battery on 08 October and 15 October
2004 are provided in Appendix D.
The clock error ranged from 10.8 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 04
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 07
Six (6) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Seven (7) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A problem was found in the new release of the CMS code. The issue is
believed to be with the namelist files, but software development is
looking into the exact cause. FDF has also been asked to provide a
User Antenna Views file, which contains SMA codes. This is needed to
complete the CMS processing of SMA supports. A meeting will be held
with FDF to discuss this issue on 21 October 2004. These two problems
have delayed the release of the new CMS code to operations. No new
date has been set for completion.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 November 2004 (DOY 307).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 October 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71582 283/1725 04-289 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71594 284/1248 04-290 PEM PEM Z/N DAY ONLY TO FULL-TIME
71595 284/1354 04-291 PEM PEM AXIS 1 & 2 ON
71607 285/0923 04-292 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71613 285/1902 04-293 MPS VT LEVEL 6
71616 286/0043 04-294 PEM (Z/H) POWER ON (OFF TO FULLTIME)
71617 286/0118 04-294 PEM (Z/H) PMON GROUP-10 DISABLED
71626 286/1607 04-295 PEM (Z/H) PEM RESET 'TIMING ERROR'
71648 288/0326 04-296 SSPP OBC MODE
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 October 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
PEHVU330VIE Y-HI HVU3 +30V CURRENT MONITOR 71629
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 October 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71582 289/1750 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71607 292/0927 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 October 2004
08 October 2004 - GMT Day 282 Beta = 55.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 -22.4/- 89.6 +4.50/+2.60 +23.2/-15.2 26.4 1.04 16.5
3 -22.4/-168.0 +3.23/+1.65 +22.0/-14.4 26.2 1.04 16.0
15 October 2004 - GMT Day 289 Beta = 64.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 -28.0/- 78.4 +3.87/+2.92 +20.0/-14.4 26.7 1.04 16.7
3 -33.6/-179.2 +2.60/+1.66 +18.8/-13.2 26.7 1.04 15.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
71613 285/1902 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 October 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-068 FOT 71572 283/0233 37190 WSGT DROPOUT/FAILOVER
04-069 FOT 71587 284/0230 37194 WSGT Equip Failover
04-070 FOT 71590 284/0559 37195 WSGT Dropout
04-071 FOT 71616 286/0043 N/A PEM IN WRONG MODE - OFF!
04-066 FOT 71632 287/0115 30832 Generic Late Acquisition
04-072 FOT 71648 288/0302 N/A SSPP In stow due to eclipse
04-066 FOT 71672 289/1743 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-067 FOT 71552 281/1730 37177 Data Drop Out
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
04-068 FOT 71572 283/0233 37190 WSGT DROPOUT/FAILOVER
04-069 FOT 71587 284/0230 37194 WSGT Equip Failover
04-070 FOT 71590 284/0559 37195 WSGT Dropout
04-071 FOT 71616 286/0043 N/A PEM IN WRONG MODE - OFF!
04-072 FOT 71648 288/0302 N/A SSPP In stow due to eclipse